# arXiv:2603.00663v1[cs.RO] 28 Feb 2026

### v1

@@ 5.2 BPRC Labeling Algorithm
We now define dominance between labels so that if label $l$ dominates label $l^{\prime}$, ${\textnormal{PartialTour}}(l)$ dominates ${\textnormal{PartialTour}}(l^{\prime})$. Consider two labels $l=(\gamma_{i,j},t,\sigma,\vec{b},\vec{g}_{\text{ub}},\vec{g}_{\text{lb}},\lambda)$ and $l^{\prime}=(\gamma_{i,j},t^{\prime},\sigma^{\prime},\vec{b}^{\prime},\vec{g}\mkern 2.0mu\vphantom{g}_{\text{ub}}^{\prime},\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime},\lambda^{\prime})$, both at target-window $\gamma_{i,j}$, and let $\Gamma={\textnormal{PartialTour}}(l)$ and $\Gamma^{\prime}={\textnormal{PartialTour}}(l^{\prime})$.
@@
    $\vec{b}[i]$ $\leq\vec{b}^{\prime}[i],\;~\forall i\in\{1,\dots,{n_{\textnormal{tar}}}\}$ (6)
@@
    $(\gamma_{i,j},\gamma_{i^{\prime},j^{\prime}})\in\mathcal{E}_{\text{tw}}\setminus\mathcal{B}$, preventing traversal of disallowed edges
@@
    $t\leq{\textnormal{LFDT}}(\gamma_{i,j},\gamma_{i^{\prime},j^{\prime}})$, ensuring that it is feasible to intercept $\gamma_{i^{\prime},j^{\prime}}$ after intercepting $\gamma_{i,j}$ at time $t$
@@
  • To compute each $\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime}[k^{\prime}]$, we first check if $t^{\prime}>\overline{t}_{i^{\prime},j^{\prime},k^{\prime}}$; if so, $\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime}[k^{\prime}]=\infty$, since it is impossible for an agent to arrive at $\xi_{i^{\prime},j^{\prime},k^{\prime}}$ by executing ${\textnormal{PartialTour}}(l^{\prime})$. If $t^{\prime}\leq\overline{t}_{i^{\prime},j^{\prime},k^{\prime}}$, we apply Bellman’s principle, similarly to (8):
@@
After generating the successor label $l^{\prime}$, if $\gamma_{i^{\prime},j^{\prime}}=\gamma_{0}$ (i.e. $l^{\prime}$ represents a tour $\Gamma^{\prime}$), we perform additional pruning checks (Lines 9-10). First, if any label $l^{\prime\prime}$ at $\gamma_{0}$ satisfies ${c^{*}_{\textnormal{red}}}({\textnormal{PartialTour}}(l^{\prime\prime}))\leq\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime}[1]-\lambda^{\prime}$, we prune $l^{\prime}$. This is because the $\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime}[1]-\lambda^{\prime}$ is a lower bound on ${c^{*}_{\textnormal{red}}}(\Gamma^{\prime})$, and if this lower bound is not smaller than the reduced cost of a previously found tour, then ${c^{*}_{\textnormal{red}}}(\Gamma^{\prime})$ cannot be smaller either. Next, if $\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime}[1]-\lambda^{\prime}\geq-\epsilon$, we prune $l^{\prime}$, since this means ${c^{*}_{\textnormal{red}}}(\Gamma^{\prime})$ is no smaller than $\epsilon$. If both of these checks fail, we compute ${c^{*}_{\textnormal{red}}}(\Gamma^{\prime})$ via an SOCP and prune $l^{\prime}$ if ${c^{*}_{\textnormal{red}}}(\Gamma^{\prime})\geq-\epsilon$. We perform the first two checks before the third to avoid the computational expense of solving an SOCP, if possible.
@@
To speed up BPRC, we apply a heuristic from (Ozbaygin et al. 2017), where each set $\mathfrak{D}[\gamma_{i,j}]$ only stores the label at $\gamma_{i,j}$ with the smallest $\min(\vec{g}_{\text{lb}})$ seen so far. This modification may prevent us from finding tours with negative reduced cost. Thus, if this modified labeling algorithm returns $\emptyset$, we then run the unmodified algorithm.
@@
8:    $l^{\prime}=(\gamma_{i^{\prime},j^{\prime}},t^{\prime},\sigma^{\prime},\vec{b}^{\prime},\vec{g}\mkern 2.0mu\vphantom{g}_{\text{ub}}^{\prime},\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime},\lambda^{\prime})=\text{SuccessorLabel}(l,\gamma_{i^{\prime},j^{\prime})})$
@@ 5.4 Theoretical Analysis
The branch-and-price paradigm finds an optimal solution to an ILP of type (1) provided that the dominance checks are correct, and the method of generating a feasible solution traversing no edge in $\mathcal{B}$, described in Section 5.3, is complete. [^7] An algorithm is complete if it generates a feasible solution when one exists and reports infeasible in finite time otherwise (Choset et al. 2005; LaValle 2006). We prove the correctness of our dominance checks here, deferring the proofs of intermediate lemmas and the proof of completeness to Appendix E.
@@
Let $T^{\prime}$ be a least-reduced-cost tour beginning with $\Gamma^{\prime}$, and let $\lambda_{\text{post}}=\sum\limits_{i=1}^{{n_{\textnormal{tar}}}}\alpha(i,T^{\prime})\lambda_{i}-\lambda^{\prime}$. Let $\tau_{\textnormal{a}}$$^{\prime}$ be a least-cost, speed-admissible trajectory executing $T^{\prime}$. Let $\tau_{\text{a,pre}}^{\prime}$ be the portion of $\tau_{\textnormal{a}}$$^{\prime}$ ending at $\gamma_{i,j}$, and let $\tau_{\text{a,post}}^{\prime}$ be the remaining portion. In Lemma 5 in Appendix E, we prove that (7) and the speed-admissibility of $\tau_{\text{a,pre}}^{\prime}$ ensure that there is a trajectory $\tau_{\text{a,pre}}$ executing $\Gamma$ satisfying
$c(\tau_{\text{a,pre}})-\lambda\leq c(\tau_{\text{a,pre}}^{\prime})-\lambda^{\prime}$ (10)
@@
Let $T^{\prime}_{\text{post}}$ be the portion of $T^{\prime}$ from $T^{\prime}[{\textnormal{Len}}(\Gamma^{\prime})+1]$ onward, and let $\sigma_{\text{post}}^{\prime}$ be its depleted capacity. Let $T$ be the concatenation of $\Gamma$ with $T^{\prime}_{\text{post}}$. Since $T^{\prime}$ is a tour, $T^{\prime}_{\text{post}}$ does not revisit targets visited by $\Gamma^{\prime}$. Combining this with (6), $T^{\prime}_{\text{post}}$ does not revisit targets visited by $\Gamma$ either, and thus $T$ does not revisit targets. Additionally, since $T^{\prime}$ is a tour, its depleted capacity $\sigma^{\prime}+\sigma_{\text{post}}^{\prime}$ is no larger than $d_{\text{max}}$; combining this with (5), the depleted capacity of $T$, i.e. $\sigma+\sigma_{\text{post}}^{\prime}$, is no larger than $d_{\text{max}}$. Therefore, since $T$ revisits no targets, satisfies capacity constraints, and begins and ends at $\gamma_{0}$, $T$ is a tour.
@@ 6 Numerical Results
We ran experiments using an Intel i9-9820X 3.3GHz CPU with 128 GB RAM, with 10 cores and hyperthreading disabled. As a baseline, we implemented the MICP (Philip et al. 2025b), with additional decision variables and constraints based on (Desrochers et al. 1987) to enforce capacity constraints. In particular, we implemented (17)-(29) from (Philip et al. 2025b), then added an additional decision variable $\sigma_{i,j}\in[0,d_{\text{max}}-d_{i}]$ for each target-window $\gamma_{i,j}$, equal to the depleted capacity immediately before visiting $\gamma_{i,j}$, if $\gamma_{i,j}$ is visited. We then added a constraint
$\sigma_{i,j}+d_{i}\leq\sigma_{i^{\prime},j^{\prime}}+(d_{\text{max}}+d_{i})(1-y_{e})$
$\;\forall e=(\gamma_{i,j},\gamma_{i^{\prime},j^{\prime}})\in\mathcal{E}_{\text{tw}}$ (15)
@@
where $y_{e}$ is a binary decision variable from (Philip et al. 2025b) which equals 1 if edge $e$ is traversed and 0 otherwise. We call this baseline Compact MICP. We also implemented two ablations of BPRC, called BPRC-ablate-dominance and BPRC-ablate-bounds. BPRC-ablate-dominance is meant to show the benefit of our dominance checks at target-windows $\gamma\neq\gamma_{0}$ by removing these checks. In particular, BPRC-ablate-dominance does not store nondominated labels at any $\gamma\neq\gamma_{0}$ and does not store $\vec{g}_{\text{ub}}$ in its labels [^8] BPRC-ablate-dominance still stores $\vec{g}_{\text{lb}}$ to break ties on the priority queue and prune labels at the depot.. Meanwhile, BPRC-ablate-bounds is meant to show the benefits of our upper and lower-bounding methods for computing $\vec{g}_{\text{ub}}$ and $\vec{g}_{\text{lb}}$, respectively. In particular, BPRC-ablate-bounds computes $\vec{g}_{\text{ub}}$ and $\vec{g}_{\text{lb}}$ exactly using SOCPs. We solved Compact MICP and all SOCPs with Gurobi 12.03.
@@
We generated problem instances by extending the instance generation method from (Bhat et al. 2024) to multiple agents. In all instances, each target had two time windows and a demand of 1. Then we varied the number of targets, the capacity, the number of agents, and the time window lengths in Experiments 1 to 4. In these experiments, we set $n_{\text{seg,tar}}$, i.e. the number of segments per target, to 32 for BPRC, 4 for BPRC-ablate-bounds, and 8 for BPRC-ablate-dominance [^9] We tuned $n_{\text{seg,tar}}$ for BPRC to achieve the best median runtime over 10 instances with 20 targets, separate from the instances in the presented experiments. BPRC-ablate-bounds and BPRC-ablate-dominance’s median runtimes were equal to the time limit for each tested $n_{\text{seg,tar}}$ value for these tuning instances, so we instead tuned the ablations on 10 instances with 15 targets.. We varied $n_{\text{seg,tar}}$ in Experiment 5. We limited each planner’s computation time to 10 min per instance. We also set a memory limit of 100 GiB, though this limit was only relevant for BPRC-ablate-dominance; the maximum measured memory usages for Compact MICP, BPRC-ablate-bounds, and BPRC were 1.41 GiB, 0.347 GiB, and 1.28 GiB, respectively.. When we report runtime statistics (e.g. median) for BPRC-ablate-dominance, we only take statistics over instances where it did not reach the memory limit.
@@
Figure 4: (a) Varying the number of targets. The gap in computation time between BPRC and the baseline and ablations widens as we increase the number of targets up to 20. The gap shrinks at 25 targets because BPRC begins reaching the time limit in many instances. We do not show a computation time for BPRC-ablate-dominance for 25 targets because it reached the memory limit in every instance. (b) Varying the capacity. BPRC shows advantage for small capacities. (c) %Gap (as defined in Experiment 2) of each method’s relaxation solution. BPRC’s relaxation has a smaller %Gap, and is thus tighter, in every instance. (d) Varying the number of agents. BPRC shows advantage when the number of agents is 3 or larger. (e) Varying the time window lengths. BPRC’s performance advantage widens as we increase time window length up to 50, then shrinks because BPRC reaches the time limit in many instances. (f) Varying the number of segments in BPRC. Horizontal and vertical axes are log-scaled. MICP’s min, median, and max are shown as flat lines without scatter points because the quantity being varied is specific to BPRC. We only show MICP’s computation time to indicate that BPRC has smaller median computation time for a wide range of parameter values.
@@
In this experiment, we varied the number of targets from 5 to 25. Each instance had 3 agents with capacity $\lceil n_{\text{tar}}/{n_{\textnormal{agt}}}\rceil$, i.e. this experiment considers instances where the agents’ capacities are severely limited. Experiment 2 considers larger capacities. The sum of time window lengths per target was set to 50. The results are shown in Fig. 4 (a). For 15 to 20 targets, BPRC’s median computation time is more than an order of magnitude smaller than Compact MICP’s. BPRC has smaller computation time than the ablations as well, indicating that our dominance checks, and our cheap methods of evaluating them, are beneficial. For 25 targets, BPRC begins reaching the time limit in some instances and thus has similar median computation time to the baseline and ablations. BPRC-ablate-dominance reached its memory limit in every instance for 25 targets, so no computation time is shown.
@@
In this experiment, we varied the capacity from 7 to 22, fixing the number of targets to 20, the number of agents to 3, and the sum of time window lengths per target to 50. The results are in Fig. 4 (b). BPRC has smaller median computation times than Compact MICP for small capacity values, and slightly larger median computation times for large capacity values. The ablations perform poorly for all capacity values.
As stated in Section 2, one reason that branch-and-price is favored over compact formulations for VRPs is that the relaxation of the integer program solved in branch-and-price tends to be tighter than the relaxations of compact formulations. To validate that this trend holds in the MT-VRP, for each instance in Experiment 2, for both BPRC and Compact MICP, we computed a lower bound $c_{\text{LB}}$ on the optimal cost by solving the convex relaxation of the method’s integer or mixed-integer program [^10] For BPRC, $c_{\text{LB}}$ is the optimal cost of LP-$\mathcal{B}$ at the root branch-and-bound node, i.e. computing it does not require enumerating $\bar{\mathcal{F}}$.. We then computed the gap of the optimal cost from each method’s lower bound, defined as $\%\text{Gap}=(c_{\text{opt}}-c_{\text{LB}})/c_{\text{LB}}*100\%.$ As shown in Fig. 4 (c), BPRC’s relaxation is empirically tighter than Compact MICP’s [^11] We only take statistics over instances where at least one method found the optimum, i.e. 8 instances for capacity 17 and 9 instances for capacity 22.. As observed for prior compact VRP formulations (Toth and Vigo 2002), Compact MICP’s median and max gap are particularly large when the capacity is small, explaining its poor performance for small capacity values.
@@
In this experiment, we varied the number of agents from 2 to 5. Each instance had 20 targets, capacity 10, and sum of time window lengths per target equal to 50. Fig. 4 (d) shows the results. BPRC’s median computation time drops steeply as we increase the number of agents from 2 to 3, then remains roughly constant as we increase the number of agents from 3 to 5. For 3 to 5 agents, BPRC’s median computation time is significantly lower than the baseline’s and the ablations’.
The large computation time for 2 agents appears to be because the agent limit (1b) is directly affecting the optimal cost when ${n_{\textnormal{agt}}}=2$, but not when ${n_{\textnormal{agt}}}>3$. That is, in 4 of the 5 instances where some method found the optimum for ${n_{\textnormal{agt}}}=2$, increasing $n_{\textnormal{agt}}$} to 3 actually resulted in a different optimal solution with a smaller cost. However, increasing $n_{\textnormal{agt}}$} from 3 to 4 did not change the optimal solution in any instance. Increasing $n_{\textnormal{agt}}$} can only decrease the optimal cost if $\lambda_{0}$ is strictly negative, which is what we often observed in the 2-agent instances: the median $\lambda_{0}$ value was -51, and the largest was -47231. Meanwhile, among all instances with ${n_{\textnormal{agt}}}>2$, the median $\lambda_{0}$ was 0, and the largest was only -889. Prior work (Afsar, Afsar, and Palacios 2021) has observed that large $\lambda_{0}$ values can slow the convergence of branch-and-price, which is what we observed here for ${n_{\textnormal{agt}}}=2$.
@@
In this experiment, we varied the sum of time window lengths per target from 10 to 70. Each instance had 20 targets and 3 agents with capacity 7. The results are in Fig. 4 (e). All methods’ computation times tend to increase with time window length. However, BPRC’s computation time is notably smaller than MICP’s for lengths 30 and 50, and smaller than the ablations’ for lengths 10, 30, and 50.
@@
In this experiment, we varied the number of segments per target, i.e. $n_{\text{seg,tar}}$ in Section 5.1, from 8 to 256, on a log-scale, i.e. we tested values of 8, 16, 32, etc. Each instance had 20 targets, 3 agents with capacity 7, and the sum of time window lengths per target equal to 50. The results are in Fig. 4 (f). BPRC’s median computation time is smaller than even the other methods’ minimum computation time for $n_{\text{seg,tar}}$ between 16 and 64 (inclusive). Thus, BPRC provides an advantage even when we change $n_{\text{seg,tar}}$ by a factor of 4.
BPRC’s computation time is large for very large $n_{\text{seg,tar}}$ because computing distances between every pair of segments becomes expensive. BPRC’s computation time is large for very small $n_{\text{seg,tar}}$ because the dominance checks become weaker, we retain more labels per target-window, and thus dominance checking becomes a bottleneck.

### v3

@@ 5.2 BPRC Labeling Algorithm
Consider two labels $l=(\gamma_{i,j},t,\sigma,\vec{b},\vec{g}_{\text{ub}},\vec{g}_{\text{lb}},\lambda)$ and $l^{\prime}=(\gamma_{i,j},t^{\prime},\sigma^{\prime},\vec{b}^{\prime},\vec{g}\mkern 2.0mu\vphantom{g}_{\text{ub}}^{\prime},\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime},\lambda^{\prime})$, both at target-window $\gamma_{i,j}$, and let $\Gamma={\textnormal{PartialTour}}(l)$ and $\Gamma^{\prime}={\textnormal{PartialTour}}(l^{\prime})$. We now define dominance so that if $l$ dominates $l^{\prime}$, ${\textnormal{PartialTour}}(l)$ dominates ${\textnormal{PartialTour}}(l^{\prime})$.
@@
    $\vec{b}[i^{\prime}]$ $\leq\vec{b}^{\prime}[i^{\prime}],\;~\forall i^{\prime}\in\{1,\dots,{n_{\textnormal{tar}}}\}$ (6)
@@
    $(\gamma_{i,j},\gamma_{i^{\prime},j^{\prime}})\in\mathcal{E}_{\text{tw}}\setminus\mathcal{B}$, preventing traversal of edges in $\mathcal{B}$
@@
    $t\leq{\textnormal{LFDT}}(\gamma_{i,j},\gamma_{i^{\prime},j^{\prime}})$, so intercepting $\gamma_{i^{\prime},j^{\prime}}$ after intercepting $\gamma_{i,j}$ at time $t$ does not violate th speed limit
@@
  • To compute each $\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime}[k^{\prime}]$, we first check if $t^{\prime}>\overline{t}_{i^{\prime},j^{\prime},k^{\prime}}$; if so, $\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime}[k^{\prime}]=\infty$, since it is impossible for an agent to intercept $\xi_{i^{\prime},j^{\prime},k^{\prime}}$ by executing ${\textnormal{PartialTour}}(l^{\prime})$. If $t^{\prime}\leq\overline{t}_{i^{\prime},j^{\prime},k^{\prime}}$, we apply Bellman’s principle, similarly to (8):
@@
After generating the successor label $l^{\prime}$, if $\gamma_{i^{\prime},j^{\prime}}=\gamma_{0}$ (i.e. $l^{\prime}$ represents a tour $\Gamma^{\prime}$), we perform additional pruning checks (Lines 9-10). First, if any label $l^{\prime\prime}$ at $\gamma_{0}$ satisfies ${c^{*}_{\textnormal{red}}}({\textnormal{PartialTour}}(l^{\prime\prime}))\leq\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime}[1]-\lambda^{\prime}$, we prune $l^{\prime}$. This is because $\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime}[1]-\lambda^{\prime}$ is a lower bound on ${c^{*}_{\textnormal{red}}}(\Gamma^{\prime})$, and if this lower bound is not smaller than the reduced cost of a previously found tour, then ${c^{*}_{\textnormal{red}}}(\Gamma^{\prime})$ cannot be smaller either. Next, if $\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime}[1]-\lambda^{\prime}\geq-\epsilon$, we prune $l^{\prime}$, since this means ${c^{*}_{\textnormal{red}}}(\Gamma^{\prime})$ is no smaller than $\epsilon$. If both of these checks fail, we compute ${c^{*}_{\textnormal{red}}}(\Gamma^{\prime})$ via an SOCP and prune $l^{\prime}$ if ${c^{*}_{\textnormal{red}}}(\Gamma^{\prime})\geq-\epsilon$. We perform the first two checks before the third to avoid the computational expense of solving an SOCP, if possible.
@@
To speed up BPRC, we apply a heuristic from (Ozbaygin et al. 2017), where each $\mathfrak{D}[\gamma_{i,j}]$ only stores the label at $\gamma_{i,j}$ with the smallest $\min(\vec{g}_{\text{lb}})$ so far. This modification may prevent finding tours with negative reduced cost. Thus, if this modified Alg. 1 returns $\emptyset$, we run the unmodified Alg. 1.
@@
8:    $l^{\prime}=(\gamma_{i^{\prime},j^{\prime}},t^{\prime},\sigma^{\prime},\vec{b}^{\prime},\vec{g}\mkern 2.0mu\vphantom{g}_{\text{ub}}^{\prime},\vec{g}\mkern 2.0mu\vphantom{g}_{\text{lb}}^{\prime},\lambda^{\prime})=\text{SuccessorLabel}(l,\gamma_{i^{\prime},j^{\prime}})$
@@ 5.4 Theoretical Analysis
Branch-and-price finds an optimal solution to an ILP of type (1) if the dominance checks are correct, and the feasible solution generation method (Section 5.3) is complete. [^9] Complete algorithms generate a feasible solution if one exists and report infeasible in finite time otherwise (Choset et al. 2005). We prove our dominance checks’ correctness here; proofs of intermediate lemmas and completeness are in Appendix E.
@@
Let $T^{\prime}$ be a least-reduced-cost tour beginning with $\Gamma^{\prime}$, and let $\lambda_{\text{post}}=\sum\limits_{i=1}^{{n_{\textnormal{tar}}}}\alpha(i,T^{\prime})\lambda_{i}-\lambda_{0}-\lambda^{\prime}$. Let $\tau_{\textnormal{a}}$$^{\prime}$ be a least-cost, speed-admissible trajectory executing $T^{\prime}$. Let $\tau_{\text{a,pre}}^{\prime}$ be the portion of $\tau_{\textnormal{a}}$$^{\prime}$ ending at $\gamma_{i,j}$, and let $\tau_{\text{a,post}}^{\prime}$ be the remaining portion. In Lemma 5 in Appendix E, we prove that (7) and the speed-admissibility of $\tau_{\text{a,pre}}^{\prime}$ ensure that there is a trajectory $\tau_{\text{a,pre}}$ executing $\Gamma$ satisfying
$c(\tau_{\text{a,pre}})-\lambda\leq c(\tau_{\text{a,pre}}^{\prime})-\lambda^{\prime}.$ (10)
@@
Let $T^{\prime}_{\text{post}}$ be the portion of $T^{\prime}$ from $T^{\prime}[{\textnormal{Len}}(\Gamma^{\prime})+1]$ onward, and let $\sigma_{\text{post}}^{\prime}$ be its depleted capacity. Let $T$ be the concatenation of $\Gamma$ with $T^{\prime}_{\text{post}}$. Since $T^{\prime}$ is a tour, $T^{\prime}_{\text{post}}$ does not revisit targets visited by $\Gamma^{\prime}$. Combining this with (6), $T^{\prime}_{\text{post}}$ does not revisit targets visited by $\Gamma$ either, and thus $T$ does not revisit targets. Also, since $T^{\prime}$ is a tour, its depleted capacity $\sigma^{\prime}+\sigma_{\text{post}}^{\prime}$ is no larger than $d_{\text{max}}$; combining this with (5), the depleted capacity of $T$, i.e. $\sigma+\sigma_{\text{post}}^{\prime}$, is no larger than $d_{\text{max}}$. Thus, since $T$ revisits no targets, satisfies capacity constraints, and begins and ends at $\gamma_{0}$, $T$ is a tour.
@@ 6 Numerical Results
We ran experiments using an Intel i9-9820X 3.3GHz CPU with 128 GB RAM, with 10 cores and hyperthreading disabled. Our baseline, called Compact MICP, is the MICP (Philip et al. 2025b), with additional decision variables and constraints based on (Desrochers et al. 1987) to enforce capacity constraints. In particular, we implemented (17)-(29) from (Philip et al. 2025b), then added an additional decision variable $\sigma_{i,j}\in[0,d_{\text{max}}-d_{i}]$ for each target-window $\gamma_{i,j}$, equal to the depleted capacity immediately before visiting $\gamma_{i,j}$, if $\gamma_{i,j}$ is visited. We then added a constraint
$\sigma_{i,j}+d_{i}\leq\sigma_{i^{\prime},j^{\prime}}+(d_{\text{max}}+d_{i})(1-y_{e})\;\forall e=(\gamma_{i,j},\gamma_{i^{\prime},j^{\prime}})\in\mathcal{E}_{\text{tw}}$ (15)
@@
where $y_{e}$ is a binary decision variable from (Philip et al. 2025b) which equals 1 if edge $e$ is traversed and 0 otherwise. We also implemented two ablations of BPRC, called BPRC-ablate-dominance and BPRC-ablate-bounds. BPRC-ablate-dominance is meant to show the benefit of our dominance checks at target-windows $\gamma\neq\gamma_{0}$ by removing these checks. In particular, BPRC-ablate-dominance does not store nondominated labels at any $\gamma\neq\gamma_{0}$ and does not store $\vec{g}_{\text{ub}}$ in its labels. [^10] BPRC-ablate-dominance still stores $\vec{g}_{\text{lb}}$ to break ties on the priority queue and prune labels at the depot. BPRC-ablate-bounds is meant to show the benefits of our upper and lower-bounding methods for computing $\vec{g}_{\text{ub}}$ and $\vec{g}_{\text{lb}}$, respectively. In particular, BPRC-ablate-bounds computes $\vec{g}_{\text{ub}}$ and $\vec{g}_{\text{lb}}$ exactly using SOCPs. We solved Compact MICP and all SOCPs with Gurobi 12.03.
@@
We generated problem instances by extending the instance generation method from (Bhat et al. 2024) to multiple agents. In all instances, each target had two time windows and a demand of 1. Then we varied the number of targets, the capacity, the number of agents, and the time window lengths in Experiments 1 to 4. In these experiments, we set the number of segments per target, i.e. $n_{\text{seg,tar}}$, to 32 for BPRC, 4 for BPRC-ablate-bounds, and 8 for BPRC-ablate-dominance. [^11] We tuned $n_{\text{seg,tar}}$ for BPRC to achieve the best median runtime over 10 instances with 20 targets, separate from the instances in the presented experiments. The ablations’ median runtimes were equal to the time limit for each tested $n_{\text{seg,tar}}$ for these tuning instances, so we instead tuned the ablations on 10 instances with 15 targets. We varied $n_{\text{seg,tar}}$ in Experiment 5. We limited each planner’s computation time to 10 min per instance. We also set a 100 GiB memory limit, which was only relevant for BPRC-ablate-dominance; the maximum measured memory usages for Compact MICP, BPRC-ablate-bounds, and BPRC were 1.41 GiB, 0.345 GiB, and 1.28 GiB, respectively. When we give runtime statistics (e.g. median) for BPRC-ablate-dominance, we only take statistics over instances where it did not reach the memory limit.
@@
Figure 4: (a) Varying the number of targets. The gap in computation time between BPRC and the baseline and ablations widens as we increase the number of targets up to 20. The gap shrinks at 25 targets because BPRC reaches the time limit in 5 instances. We do not show a computation time for BPRC-ablate-dominance for 25 targets because it reached the memory limit in every instance. (b) Varying the capacity. BPRC shows advantage for small capacities. (c) %Gap (as defined in Experiment 2) of each method’s relaxation solution. BPRC’s relaxation has a smaller %Gap, and is thus tighter, in every instance. (d) Varying the number of agents. BPRC shows advantage when the number of agents is 3 or larger. (e) Varying the time window lengths. BPRC’s performance advantage widens as we increase time window length up to 50, then shrinks because BPRC reaches the time limit in many instances. (f) Varying the number of segments in BPRC. Horizontal and vertical axes are log-scaled. MICP’s min, median, and max are shown as flat lines without scatter points because the quantity being varied is specific to BPRC. We only show MICP’s computation time to indicate that BPRC has smaller median computation time for a wide range of parameter values.
@@
We varied $n_{\textnormal{tar}}$} from 5 to 25. Each instance had 3 agents with capacity $\lceil n_{\text{tar}}/{n_{\textnormal{agt}}}\rceil$, i.e. this experiment considers instances where capacity is severely limited. Experiment 2 considers larger capacities. The sum of time window lengths per target was 50. The results are in Fig. 4 (a). For 15 to 20 targets, BPRC’s median runtime is more than an order of magnitude smaller than Compact MICP’s. BPRC has smaller runtime than the ablations as well, indicating that our dominance checks, and our cheap methods of evaluating them, are beneficial. For 25 targets, BPRC reaches the time limit in 5 instances and thus has similar median runtime to the baseline and ablations. For 25 targets, pricing was the bottleneck, taking more than 80% of the runtime in 8 of the 10 instances.
@@
We varied the capacity from 7 to 22, fixing the number of targets to 20, the number of agents to 3, and the sum of time window lengths per target to 50. The results are in Fig. 4 (b). BPRC has smaller median runtimes than Compact MICP for small capacities, and slightly larger median runtimes for large capacities. The ablations perform poorly for all capacities.
As stated in Section 2, one reason that branch-and-price is favored over compact formulations for VRPs is that the relaxation of branch-and-price’s integer program tends to be tighter than the relaxations of compact formulations. To validate that this holds in the MT-VRP, for each instance in Experiment 2, for both BPRC and Compact MICP, we computed a lower bound $c_{\text{LB}}$ on the optimal cost by solving the convex relaxation of the method’s integer or mixed-integer program. [^12] For BPRC, $c_{\text{LB}}$ is the optimal cost of LP-$\mathcal{B}$ at the root branch-and-bound node, i.e. computing it does not require enumerating $\mathcal{S}$. We then computed the gap of the optimal cost from each method’s lower bound, defined as $\%\text{Gap}=(c_{\text{opt}}-c_{\text{LB}})/c_{\text{LB}}*100\%.$ As shown in Fig. 4 (c), BPRC’s relaxation is empirically tighter than Compact MICP’s. [^13] We only take statistics over instances where at least one method found the optimum, i.e. 8 instances for capacity 17 and 9 instances for capacity 22.
@@
We varied the number of agents from 2 to 5. Each instance had 20 targets, capacity 10, and sum of time window lengths per target equal to 50. Fig. 4 (d) shows the results. BPRC’s median computation time drops steeply as we increase the number of agents from 2 to 3, then remains roughly constant as we increase the number of agents from 3 to 5. For 3 to 5 agents, BPRC’s median computation time is significantly lower than the baseline’s and the ablations’.
The large runtime for 2 agents appears to be because the agent limit (1b) is directly affecting the optimal cost when ${n_{\textnormal{agt}}}=2$, but not when ${n_{\textnormal{agt}}}>3$. That is, in 4 of the 5 instances where some method found the optimum for ${n_{\textnormal{agt}}}=2$, increasing $n_{\textnormal{agt}}$} to 3 actually resulted in a different optimal solution with a smaller cost. However, increasing $n_{\textnormal{agt}}$} from 3 to 4 did not change the optimal solution in any instance. Increasing $n_{\textnormal{agt}}$} can only decrease the optimal cost if $\lambda_{0}$ is strictly negative, which is what we often observed in the 2-agent instances: the median $\lambda_{0}$ value was -52, and the largest was -47231. Meanwhile, among all instances with ${n_{\textnormal{agt}}}>2$, the median $\lambda_{0}$ was 0, and the largest was only -889. Prior work (Afsar, Afsar, and Palacios 2021) has observed that large $\lambda_{0}$ values can slow the convergence of branch-and-price, which is what we observed here for ${n_{\textnormal{agt}}}=2$.
@@
We varied the sum of time window lengths per target from 10 to 70. Each instance had 20 targets and 3 agents with capacity 7. The results are in Fig. 4 (e). All methods’ runtimes tend to increase with time window length. BPRC’s median runtime is smaller than Compact MICP’s for lengths 30 and 50, and smaller than the ablations’ for lengths 10, 30, and 50.
@@
We varied $n_{\text{seg,tar}}$ from 8 to 256, on a log-scale, i.e. we tested values 8, 16, 32, etc. Each instance had 20 targets, 3 agents with capacity 7, and sum of time window lengths per target equal to 50. The results are in Fig. 4 (f). BPRC’s median runtime is smaller than the other methods’ minimum runtime for $n_{\text{seg,tar}}$ from 16 to 64 (inclusive). Thus, BPRC shows advantage even when we change $n_{\text{seg,tar}}$ by a factor of 4.
BPRC’s runtime is large for large $n_{\text{seg,tar}}$ because computing distances between all pairs of segments becomes expensive. BPRC’s runtime is large for small $n_{\text{seg,tar}}$ because dominance checks become weaker, causing more retained labels per target-window, making pricing expensive.